OpenStudy (anonymous):

The slope of a line is 2, and the y-intercept is 0. What is the equation of the line written in slope-intercept form?

OpenStudy (anonymous):

a. y = x + 2 b. y = 2 c. y = 2x

OpenStudy (anonymous):

y = mx + b where m is the slope of the line and b is the y-intercept of the line, or the y-coordinate of the point at which the line crosses the y-axis. To write an equation in slope-intercept form, given a graph of that equation, pick two points on the line and use them to find the slope. This is the value of m in the equation. Next, find the coordinates of the y -intercept--this should be of the form (0, b) . The y - coordinate is the value of b in the equation. Finally, write the equation, substituting numerical values in for m and b . Check your equation by picking a point on the line (not the y -intercept) and plugging it in to see if it satisfies the equation.

OpenStudy (anonymous):

y=2x+0